# PreferThinker: Reasoning-based Personalized Image Preference Assessment

### v2

@@ 4.3 Ablation Study and Discussion
Table 2: Ablation study of PreferThinker on PreferImg. Ass.
: Assessment accuracy. Pred.: Profile prediction accuracy.
Base SFT RL PR Seen-SP Unseen-SP Seen-MP Unseen-MP Ass. Pred. Ass. Pred. Ass. Pred. Ass. Pred. 75.4 70.4 72.0 70.88 62.0 71.1 64.8 71.1 92.0 84.2 91.8 85.3 81.2 73.9 81.6 74.2 89.6 79.6 91.0 80.3 78.4 76.9 77.2 73.5 94.9 84.1 95.2 85.2 90.4 72.5 89.2 74.1 96.6 85.4 96.4 86.2 92.0 78.6 92.8 80.7
@@ Appendix A Experimental Settings
We utilize Qwen2.5-VL-7B Instruct model as our base model and perform supervised fine-tuning using LLaMA-Factory (zheng2024llamafactory). We train the model on the proposed CoT-style cold-start dataset PreferImg-CoT for one epoch. The dataset includes 50,000 user samples with a single preference and 10,000 user samples with multiple preferences. we freeze the parameters of the vision tower and the multi-modal projector, and fine-tune the LLM. We employ DeepSpeed Zero-3 optimization strategy to handle the memory requirements of large models. The training settings of SFT are detailed in the Table S2.
@@ Appendix B Dataset Details
B.1 Visual Preference Elements.
To ensure the diversity of personalized visual preference profiles, we collect an extensive set of vocabulary related to five key visual elements from Lexica’s  [^2] https://lexica.art/, as illustrated in Fig. B.1. Leveraging this lexicon, we construct a large-scale and rich set of visual preference profiles, laying a robust foundation for constructing the dataset PreferImg and CoT-style dataset PreferImg-CoT.
@@ D.1 Visualization Results on Users with Single Preference
We present more visualization results on the PreferImg where users’s reference images reflect single preference in Figures S8, S9 and S10. It is observed that PreferThinker first predicts the users’ preference and non-preference profiles from the reference images, and then provides a reliable assessment and score for candidate images based on these profiles, which not only enables accurate assessment but also provides interpretability.
@@ D.2 Visualization Results on Users with Multiple Preferences
We show more visualization results on the PreferImg where users’s reference images reflect multiple Preferences in Fig. S11 and S12. Our model can accurately predict multiple preference profiles from reference images and provide a reliable assessment and score for candidate images based on these profiles, revealing the effectiveness of PreferThinker for users with multiple preferences.
@@ D.3 Visualization Results on Real Users
To demonstrate the generalizability of our method to real-world users, we further utilize the PreferThinker to conduct personalized preference assessment for real users from PickaPic. As shown in Figures S13, S14, S15 and S16 Although the preferences in real users’ reference images are more complex, our method can still effectively extract their primary preference profiles and provide a correct assessment.
@@ E.4 Limitation and Future work
The proposed PreferThinker has shown promising results on The PreferImg benchmark. While the PreferImg dataset features diverse personalized visual preferences and large-scale user samples, its manual character of personalized preferences still creates a gap between it and the complex personalized preferences of real users. This may cause the model to make inaccurate user profile predictions when handling real users with complex preferences, which in turn leads to flawed reasoning and incorrect assessments, as shown in Fig. S21. We will further explore more complex personalized preferences and real user data to enhance the model’s generalizability to real-world scenarios in future work.

### v3

@@ 4.3 Ablation Study and Discussion
Table 2: Ablation study of PreferThinker. Ass.: Assessment accuracy. Pred.: Profile prediction accuracy.
Base SFT RL PR Seen-SP Unseen-SP Seen-MP Unseen-MP
Ass. Pred. Ass. Pred. Ass. Pred. Ass. Pred.
75.4 70.4 72.0 70.88 62.0 71.1 64.8 71.1
92.0 84.2 91.8 85.3 81.2 73.9 81.6 74.2
89.6 79.6 91.0 80.3 78.4 76.9 77.2 73.5
94.9 84.1 95.2 85.2 90.4 72.5 89.2 74.1
96.6 85.4 96.4 86.2 92.0 78.6 92.8 80.7
Table 3: Performance and speed of PreferThinker without explicit CoT.
@@
PreferImg PickaPic
ViPer 86.0 62.2 2.19 s
OpenAI-GPT-4o 88.0 65.7 12.92 s
Claude 3.7 88.3 64.9 18.22 s
Doubao-1.5-vision 84.3 63.8 5.39 s
Ours (w/o CoT) 93.2 65.2 0.80 s
@@ 5 Conclusion
Acknowledgments
This work was supported by National Key RD Program of China under Grant No. 2022YFA1004100.
@@ Appendix A Experimental Settings
We utilize Qwen2.5-VL-7B Instruct model as our base model and perform supervised fine-tuning using LLaMA-Factory (Zheng et al., 2024). We train the model on the proposed CoT-style cold-start dataset PreferImg-CoT for one epoch. The dataset includes 50,000 user samples with a single preference and 10,000 user samples with multiple preferences. we freeze the parameters of the vision tower and the multi-modal projector, and fine-tune the LLM. We employ DeepSpeed Zero-3 optimization strategy to handle the memory requirements of large models. The training settings of SFT are detailed in the Table S2.
@@ Appendix B Dataset Details
B.1 Visual Preference Elements of Preference Profile.
To formalize a user’s complex personalized visual preferences, we identify 15 visual elements that most frequently appear in Lexica’s  [^2] https://lexica.art/ text prompts and strongly influence user preference toward personalized image generation. We then conduct a user study with 100 participants, asking each to select the five most important visual elements as visual preference profile. The result in Fig S1(a) reveals that art style, color, detail, art medium and saturation are voted as the most representative visual elements for characterizing personalized preferences. To ensure preference profile diversity, we collect a rich vocabulary of related terms for each visual element, as shown in Fig. S1(b), laying a solid foundation for constructing a large-scale and diverse dataset for personalized image preference assessment.
To ensure the diversity of personalized visual preference profiles, we collect an extensive set of vocabulary related to five key visual elements, as illustrated in Fig. S2. Leveraging this lexicon, we construct a large-scale and rich set of visual preference profiles, laying a robust foundation for constructing the dataset PreferImg and CoT-style dataset PreferImg-CoT.
@@
Figure S1: Key visual elements of preference profile. (a) User study result reveals that color, art style, art medium, saturation, and detail are voted five key elements representing the visual preference profile. (b) World clouds show that each element has a rich vocabulary associated with it.
@@ D.1 Visualization Results on Users with Single Preference
We present more visualization results on the PreferImg where users’s reference images reflect single preference in Figures S9, S10 and S11. It is observed that PreferThinker first predicts the users’ preference and non-preference profiles from the reference images, and then provides a reliable assessment and score for candidate images based on these profiles, which not only enables accurate assessment but also provides interpretability.
@@ D.2 Visualization Results on Users with Multiple Preferences
We show more visualization results on the PreferImg where users’s reference images reflect multiple Preferences in Fig. S12 and S13. Our model can accurately predict multiple preference profiles from reference images and provide a reliable assessment and score for candidate images based on these profiles, revealing the effectiveness of PreferThinker for users with multiple preferences.
@@ D.3 Visualization Results on Real Users
To demonstrate the generalizability of our method to real-world users, we further utilize the PreferThinker to conduct personalized preference assessment for real users from PickaPic. As shown in Figures S14, S15, S16 and S17 Although the preferences in real users’ reference images are more complex, our method can still effectively extract their primary preference profiles and provide a correct assessment.
@@ E.4 Limitation and Future work
The main limitation of our method is that the personalized preference dataset built on synthetic profiles may not fully capture the complex and subtle real human preferences. Our current preference profile only contains five key visual attributes that influence human visual preferences. However, real users’ preferences are often complex and subtle, shaped by multiple factors. Beyond visual attributes, they are also closely related to image content semantics and users’ backgrounds (e.g., age, gender and culture). Consequently, our method may struggle with complex and hard-to-articulate real user preferences, as shown in Fig. S22. In future work, we will incorporate additional key factors into preference profiles to more comprehensively capture real users’ preferences.